1. Nitric acid (HNO3) is a strong acid that is completely ionized in aqueous solutions of concentrations ranging from 1% to 10% (1.50 M ). However, in more concentrated solutions, part of the nitric acid is present as un-ionized molecules of HNO3. For example, in a 50% solution (7.50 M ) at 25°C, only 33% of the molecules of HNO3 dissociate into H+ and NO3–. What is the value of Ka for HNO3?
1
Expert's answer
2020-04-13T02:24:52-0400

We use the Ostwald breeding law:


"K_a = {C\alpha ^2 \over 1 - \alpha}"

"K_a = {7.50*(0.33)^2 \over 1 - 0.33} = 1.22"

Answer: 1.22